OpenStudy (anonymous):

Solve the following proportion: s/7 = 27/21

OpenStudy (anonymous):

you will get a metal if you help

OpenStudy (anonymous):

Let's solve your equation step-by-step. s7=2721 Step 1: Cross-multiply. s7=2721 s*(21)=(27)*(7) 21s=189 Step 2: Divide both sides by 21. 21s21=18921 s=9 Answer: s=9

OpenStudy (anonymous):

thank you @aloud

OpenStudy (anonymous):

anytime